Write the equation of the line passing through (3, –1) and (3, 6).
Found 2 solutions by stanbon, tutorcecilia:
Answer by stanbon(75887)

Write the equation of the line passing through (3, –1) and (3, 6).
-------------
Find the slope: m = [6--1]/[3-3] = undefined because the denominator is zero
So the slope is undefined; the line is a vertical line.
------------
Since the line passes thru two points with x=3
the equation of the line is x=3.
===============
Cheers,
Stan H.

Answer by tutorcecilia(2152)

(3, –1) and (3, 6)
Use the formula for the slope of a line:
Slope = m = (y1-y2)/(x1-x2)
Plug-in the values and solve: m = (-1-6)/(3-3) = -7/0=undefined
Since the slope is undefined, the line is a straight vertical line passing through only one point on the x-axis. Since all x-values are 3, the equation is:
x=3